OpenStudy (anonymous):

A 250 uF capacitor is connected in series with a 15ohm resistor to a 230V 50Hz supply calculate Impedance (Z)

OpenStudy (anonymous):

I managed to get Xc to be 12.73 but cant find inductance for Xl

OpenStudy (anonymous):

XL can be calculated using the equation XL= 2pfL. f= frequency in hertz L= inductance in henries.

OpenStudy (anonymous):

L= x/ (2(pi)(f))

OpenStudy (anonymous):

I understand so far but this is the part I get stuck on, They haven't given inductive reactance so I don't think you can use L= x/ (2(pi)(f)) On my formula sheet it says L i=Vs/R which I tried and got 15.33 So Xl=4816 ohms but that seems a bit big

OpenStudy (anonymous):

I got something much smaller than that. I think you should just use L= x/ (2(pi)(f)) for the sake of getting the right answer

OpenStudy (anonymous):

Could you show me how you did it please?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

cheers